OpenStudy (anonymous):

Two numbers, x and y, are randomly chosen from the interval [0,1]. What is the expected value of (x+y)?

OpenStudy (kainui):

Without doing any math, what do you expect the value to be? Just like, make a guess.

OpenStudy (anonymous):

just what i was going to say before you do any math, make a hypothesis there is only one answer that should make any sense

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

i know its about 1/2 but i don't know the exact answer

OpenStudy (kainui):

What is the maximum and minimum value of (x+y) ?

OpenStudy (anonymous):

it is one half for one of the random variables, not for the sum of the two

OpenStudy (anonymous):

is it 0.75

OpenStudy (kainui):

What is the maximum and minimum value of (x+y) ?

OpenStudy (anonymous):

1,0

OpenStudy (kainui):

Can't we pick x=1 and y=1 or x=0 and y=0? This would give us (1+1) and (0+0) for our max and min, meaning we have 2 and 0 as the range of possible values.

OpenStudy (anonymous):

@Kainui maybe it is time to "do the math"

OpenStudy (kainui):

Ok, I'll let you do the math if that's what you want @satellite73 ;P

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no i don't want to really but i thought the answer would be obvious

OpenStudy (anonymous):

\[E(x+y)=E(x)+E(y)\] here and since these are identical random variables this is the same as \(2E(x)\)

OpenStudy (anonymous):

\[E(x)=\int_{-\infty}^{\infty}xP(x)\] in this case \[P(x)= \left\{\begin{array}{rcc} 1 & \text{if} &0<x<1 \\ 0& \text{otherwise} \end{array} \right. \]

OpenStudy (anonymous):

so \[E(x)=\int _0^1xdx=\frac{1}{2}\]

OpenStudy (anonymous):

so 1/2

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\(E(x)=\frac{1}{2}. 2E(x)=1\)

OpenStudy (anonymous):

really it should be way more obvious than all this silly math they are uniform the smallest is zero the largest is two and the average is one

OpenStudy (kainui):

Let's pretend you have two friends. Each of them are either going to give you $0 or $1. So what amount of money could you possibly get? Well they could either both give you 1 dollar, or both 0 dollars, or they could give you either 1 or 0 dollars. 1+1=2 0+0=0 1+0=1 0+1=1 It looks like out of all 4 possibilities one of these is more likely than the others because it appears twice. Why is that?

OpenStudy (anonymous):

Since we can add expected values of independent events, the expected value of the sum is the sum of the expected values of each. Since we are just as likely to pick x as 1-x, the expected value of x (and likewise y) is equal to \frac12. Therefore the expected value of the sum is \frac12+\frac12=\boxed1.

OpenStudy (anonymous):

wait i just figure it out

OpenStudy (anonymous):

these are dependent events. Sry for not realizing fast enough

OpenStudy (kainui):

Wait... They are independent according to what you told us in the question. Whatever x is has no bearing on what y is and vice versa.

OpenStudy (anonymous):

0<=x+y<=2

OpenStudy (anonymous):

sry typo